I need help factoring x^2+x-30
Answer by uma(370)

x^2 + x - 30.
For factorising this expression, we need to find two integers whose product is -30 and whose sum is 1.
On observation we find the integers to be 6, -5
So x^2 + x - 30
= x^2 + 6x - 5x - 30 [splitting the middle term]
= x(x + 6) - 5(x + 6) [removing the common factor]
= (x - 5) (x + 6)
Thus x^2 + x - 30 = (x-5)(x+6).
Good Luck!!!